11. 1,000 mL of NS must infuse via IV over 8 hours. The drop factor is 20 gtt/mL. Find the flow rate in gtt/min.
Answer #1

ANSWER.

11) Calculation of flow rate.

           Flow rate(gtt/min) = Volume(ml) / Time(min) * Drop factor(gtt/ml)

     Here

         Volume(ml) = 1000 ml

          Time(min) = 480 min (8hr)

          Drop factor(gtt/ml) = 20 gtt/ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                       = 1000/480 * 20

                                                       = 2.08 * 20

                                                       = 41.6 gtt/min.